How many photons per second are emitted by a 7.50π‘šπ‘Š 𝐢𝑂2 laser that has a wavelength of 10.6πœ‡π‘š?


1
Expert's answer
2021-12-17T18:16:32-0500

"P=\dot nh\frac c{\lambda},"

"\dot n=\frac{P\lambda}{hc}=4\cdot 10^{15}~s^{-1}."